OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to x from the first equation is substituted into the second equation. x – 2y = 2 x + 3y = 17 x + 3(2y + 2) = 17 x + 3(–2y + 2) = 17 (2y + 2) + 3y = 17 (–2y + 2) + 3y = 17 If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to x from the first equation is substituted into the second equation. x – 2y = 2 x + 3y = 17 x + 3(2y + 2) = 17 x + 3(–2y + 2) = 17 (2y + 2) + 3y = 17 (–2y + 2) + 3y = 17 @Mathematics

OpenStudy (tcp2315):

first rearrange the first equation for x.. x=2y + 2 then plug it into the second (2y+2) +3y =17 3rd answer

OpenStudy (anonymous):

x = 2 + 2y 2+2y + 3y = 17 5y+2 = 17 5y = 15 y = 3

OpenStudy (anonymous):

oh then it's the 3rd one